OpenStudy (anonymous):

I will fan and medal!! What are the explicit equation and domain for an arithmetic sequence with a first term of 8 and a second term of 5?

OpenStudy (anonymous):

A. an=8-5(n-1) ; all integers where n is greater than or equal to 1 B. an= 8-5(n-1) ; all integers where n is greater than or equal to 0 C. an=8-3(n-1) ; all integers where n is greater than or equal to 1 D. an=8-3(n-1) ; all integers where n is greater than or equal to 0

OpenStudy (anonymous):

Also this one What is the sum of the geometric sequence -3, 18, -108 if there are 8 terms? A. 719,835 B. 119,973 C. -119,973 D. -719,835